A sledge moving over a smooth horizontal surface of ice at a velocity v2 drives out on a horizontal road and comes to a halt. The sledge has a length l, mass m and friction between runners and road is mhu

Grade:12

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Please tell what is to be done in question

Regards
Khimraj
4 years ago
Assume that the spring compression is x .
Here , the Initial kinetic energy of the block 
K.E = 1/2 mv²
When the spring is compressed at it maximum range at that point the Kinetic energy will be changed to potential energy .
P.E = 1/2 mx²
Now from the law of conservation of energy .we get
1/2 mv² = 1/2 kx²
mv² = kx²
v² = kx²/m
x² = (v² × m) /k
The maximum compression x, is 
⇒Now for the second case , we assume that there is no loss in the spring. So the kinetic energy is the same when spring comes to it's original position.
But velocity is a vector quantity , so it will not be same as like original . 

Hope it Helps.